### v1

# In-Context Multi-Operator Learning with DeepOSets

Abstract
In-context Learning (ICL) is the remarkable capability displayed by some machine learning models to learn from examples in a prompt, without any further weight updates. ICL had originally been thought to emerge from the self-attention mechanism in autoregressive transformer architectures. DeepOSets is a non-autoregressive, non-attention based neural architecture that combines set learning via the DeepSets architecture with operator learning via Deep Operator Networks (DeepONets). In a previous study, DeepOSets was shown to display ICL capabilities in supervised learning problems. In this paper, we show that the DeepOSets architecture, with the appropriate modifications, is a multi-operator in-context learner that can recover the solution operator of a new PDE, not seen during training, from example pairs of parameter and solution placed in a user prompt, without any weight updates. Furthermore, we show that DeepOSets is a universal uniform approximator over a class of continuous operators, which we believe is the first result of its kind in the literature of scientific machine learning. This means that a single DeepOSets architecture exists that approximates in-context any continuous operator in the class to any fixed desired degree accuracy, given an appropriate number of examples in the prompt. Experiments with Poisson and reaction-diffusion forward and inverse boundary-value problems demonstrate the ability of the proposed model to use in-context examples to predict accurately the solutions corresponding to parameter queries for PDEs not seen during training.
1 Introduction
In learning the solution operator of a parametric PDE from sample data, one is not interested in merely obtaining a single solution for a single PDE, but to learn the mapping between parameter and solution, so that new solutions corresponding to new values of the parameter can be obtained quickly and without further training. Furthermore, both the parameter and the solution are typically functions in infinite-dimensional spaces. For example, the problem might be to learn the space-varying fluid pressure (solution) corresponding to the space-varying permeability in a porous medium (parameter), where the physical process is governed by Darcy’s Equation. The training data in this case consists of pairs of parameter and corresponding solution, obtained from experiments or, more commonly, by solving Darcy’s equation using a high-fidelity solver. The trained operator allows quick exploration of the parameter space, which is unfeasible to do with expensive experiments or high-fidelity solves. The most well-known operator learners are DeepONets ludeeponet2021 and Fourier Neural Operators li2020fourier, which are both based on specific deep neural network architectures.
A more recent development is foundation models for scientific problems, which is inspired by the success of of large language models. A foundation model can be trained on a large body of data and be adapted for new tasks, perhaps with no further training bommasani2021opportunities. In the context of PDEs, foundation models correspond to multiple operator (multi-operator) learning. The challenge could be to be able to train a model with data from diverse PDEs with diverse equation coefficients and boundary conditions, and then produce fast solutions for new PDEs on demand. In contrast to single operator learning, here a well-posedness question arises, since multiple solution operators acting on the same parameter function will produce different solutions. How is the model supposed to identify the correct operator? There are two basic methods to resolve this issue: the first, is to fine-tune the model on a small amount of data from the new PDE; the second is to use in-context learning (ICL), when the PDE is identified from example pairs in a prompt with no further training.
In-Context Operator Networks (ICON) yang2023context and its subsequent variants yang2024pde; yang2023ICON-LM; zhang2025GenICON remain the only published approach to in-context learning of PDE solution operators (see Section 2 for other work related to multi-operator learning and foundation models for PDEs). ICON is a framework that adapts GPT-style in-context learning to PDE problems, by employing a decoder transformer to generate the solution from a sequential prompt followed by the query. ICON is thus an autoregressive, attention-based approach, with quadratic complexity in prompt length, due to the dot-product self-attention mechanism. This can introduce significant computational overhead and loss of accuracy with very long prompts. In addition, if the query changes, the prompt has to be processed sequentially again from the beginning (unless caching is used).
In this paper, we propose an alternative approach to in-context learning of PDE solution operators, which processes the prompt non-autoregressively, in a fully parallel manner, without using self-attention, being therefore much faster during both training and inference time. This approach is based on DeepOSets, a neural architecture that combines set learning via the DeepSets architecture zaheer2017deep with operator learning via the previously mentioned DeepONets architecture. In DeepOSets, the DeepSets layer acts as an encoder of the prompt into a latent representation that is the input, along with the query, to the branch path of a DeepONet. The DeepSets layer confers linear complexity in the length of the prompt, in contrast to the quadratic complexity of self-attention, and a built-in permutation-invariance inductive bias to the order of the examples in the prompt.
DeepOSets was shown in a previous publication to display ICL capabilities in regression problems chiu2024deeposets. Here, we introduce architectural modifications to DeepOSets that allow it to act as a multi-operator learner for parametric PDEs. This is accomplished by discretizing a variable number of solution and parameter function pairs, which are entered into the prompt and encoded by the DeepSets layer into latent representation. The latter is concatenated to the discretization of the query parameter function and entered directly into the branch path of the DeepONet.
Crucially, we prove that this DeepOSets architecture approximates uniformly a compact set of continuous solution operators, which means that DeepOSets are not only universal approximators of single solution operators, like DeepONets, but they are universal approximators of multiple operators, in the sense that a single DeepOSets architecture exists that can approximate any operator in the class to the same fixed degree of accuracy. This guarantees that DeepOSets can learn in-context the solution operators of multiple PDEs.
Experiments with Poisson and reaction-diffusion forward and inverse problems demonstrate that DeepOSets can accurately predict the solution to a new PDE with inverse or forward problem and unknown value of equation coefficients and boundary conditions learned in-context.
2 Related Work
MIONet is probably the first approach to attempt multi-input operator learning, by encoding several input functions via multiple branch network (and a shared trunk network) of a DeepONet, though it still learns a single operator jin2022mionet. POSEIDON is a large foundation model for PDEs based on a scalable Operator Transformer (scOT) with a vision transformer architecture, is pretrained on compressible Euler and incompressible Navier-Stokes equations and applied to 15 downstream tasks, showing better sample efficiency and accuracy than baselines herde2024poseidon. PROSE-PDE is a transformer-based multi-operator learning approach trained on one-dimensional time-dependent nonlinear PDEs that encodes both numerical and symbolic data types for forward and inverse PDE problems sun2025towards. LeMON is based on pretraining and fine-tuning for multi-operator learning, and introduces low-rank adaptation to reduce computational cost sun2024lemon. MT-DeepONet introduces modifications to the branch network of a DeepONet to accommodate various functional forms of parameterized coefficients and multiple geometries kumar2025synergistic. In-context learning for PDE solution operators was introduced by ICON yang2023context. The follow-up study on 1D scalar conservation laws provides a detailed methodology for applying ICON to temporally evolving PDEs yang2024pde, while ICON-LM integrates ICON concepts with language-model architectures yang2023ICON-LM. GenICON introduces a generative, probabilistic approach to ICON zhang2025GenICON.
3 The DeepOSets Architecture for In-Context Operator Learning
The DeepOSets model is a hybrid neural network comprising a DeepSets encoder and a DeepONet decoder, as seen in Figure 1. The DeepSets module encodes non-autoregressively the discretized parameter and solution function example pairs in the prompt. Each pair $(u_{i},G(u_{i}))$ is first mapped to an individual fixed-length embedding $\bf{h}_{i}$ by a shared multi-layer perceptron (MLP) denoted $\phi$. These intermediate representations are then combined by a pooling layer into the embedding vector $\bf{h}$. This introduces a permutation invariance inductive bias to the prompt examples, as long as the pooling layer is a permutation invariant operation, such as min, max, mean or set attention pooling. The DeepONet takes the concatenation of the embedding $\bf{h}$ and the discretized query parameter function $u_{q}$ at its branch path and the query coordinates $x_{q}$ at its trunk path, and outputs the predicted solution function $\hat{G}(u_{q})$ evaluated at $x_{q}$. The output is therefore entirely mesh-free.
Figure 1: DeepOSets architecture for in-context multi-operator learning.
4 Universality of DeepOSets for In-Context Operator Learning
In this section, we prove that DeepOSets are universal uniform approximators over a compact set of continuous operators, i.e., a single DeepOSets neural network exists that bounds the supremum of the error over the entire class to any desired fixed accuracy. This shows that a unique DeepOSets neural network could be learned that approximates in-context any continuous operator to the desired accuracy, given an appropriate number of examples in the prompt.
We first define the class of operators of interest. Let $X:=C(K_{1}),~Y:=C(K_{2})$, with $K_{1},K_{2}\subset\mathbb{R}^{d}$ compact subsets of some Euclidean space. Let $V\subset X$ be compact. Given this, consider the space of continuous operators (in the induced operator norm):
$\mathcal{O}:=\{G:V\rightarrow Y~\text{continuous}\}$
with norm given by
$\|G\|_{\mathcal{O}}:=\sup_{x\in V}{\|G(x)\|_{Y}}.$
We let $\mathcal{K}\subset\mathcal{O}$ to be compact subset of this space of operators. We will need to discretize functions in $X$ on different grids in $K_{1}$ and will also need to consider certain coverings of the set $V$. For this purpose, we introduce the following definition.
Definition 4.1 (Grid).
Let $\delta>0$ and $C>1$. A $(\delta,C)$-discretization of a compact metric space $K$ is a set of points $\textbf{x}=\{x_{1},\cdots,x_{n}\}$ such that
$|\{x_{i}\in B(x,\delta/2)\}|<C|\{x_{i}\in B(y,\delta)\}|$
for any $x,y\in K$.
Intuitively, a $(\delta,C)$-discretization is a $\delta$-net of $K$ which is not too unbalanced, in the sense that the density of points is relatively uniform. Let us first note that this definition is not vacuous.
Lemma 4.1.
Let $K$ be a compact set in a metric space, $\delta>0$ and $C>1$. Then there always exists a $(\delta,C)$-discretization of $K$.
Proof.
See the Appendix.∎
Given a continuous function $u:K\rightarrow\mathbb{R}$ and a $(\delta,C)$-discretization $\textbf{x}=(x_{1},...,x_{n})$, we denote by
$u(\textbf{x}):=\left(u(x_{i})\right)_{i=1}^{n}\in\mathbb{R}^{n}$ (1)
for the function evaluated at the grid x. Given these definitions and notation, our main result is the following.
Theorem 4.2 (Uniform universality for In-Context Operator Learning).
Given $\varepsilon>0$, $C>1$ and a compact set $\mathcal{K}\subset\mathcal{O}$, there exist integers $N,M\geq 1$, $\delta>0$, fixed discretizations $\textbf{x}=\{x_{1},...,x_{k}\}$ of $K_{1}$ and $\{y_{1},...,y_{n}\}$ of $K_{2}$, and continuous functions $\Psi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{M}$, $\rho:\mathbb{R}^{k+N}\rightarrow\mathbb{R}^{M}$, and $\Phi:\mathbb{R}^{k+n}\rightarrow\mathbb{R}^{N}$, such that
$\sup_{G\in\mathcal{K}}\>\left|\,G(u_{q})(y)-\Psi(y)\cdot\rho\left(u_{q}(\textbf{x}),\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))\right)\right|<\varepsilon.$
for any $(\delta,C)$-discretization $\textbf{u}:=(u_{1},\cdots,u_{m})$ of $V$, any $u_{q}\in V$, and any $y\in K_{2}$.
Proof.
See the Appendix.∎
The continuous functions guaranteed by Theorem 4.2 can be approximated using neural networks by invoking standard universal approximation theorems for MLPs cybenko1989approximation; hornik1991approximation; pinkus1999approximation. Here it is important that we are averaging the outputs of $\Phi$ to ensure that the function $\rho$ only needs to be approximated on a compact set.
5 Experimental Results
Here, we demonstrate that the proposed DeepOSets architecture can learn the solution operator for multiple PDEs, with the correct PDE being disambiguated with the prompt examples, without any further weight updates. We consider two one-dimensional parametric PDEs and forward and inverse problems for each.
The Poisson boundary-value PDE reads:
$\frac{d^{2}u(x)}{dx^{2}}$ $=\,f(x)\,,\quad 0<x<1\,$, (2)
with $u(0)=u_{0}$ and $u(1)=u_{1}$. The forward problem is to find the solution $u(x)$ given the forcing function $f(x)$. The inverse or control problem is to find the forcing function $f(x)$ that produces the given trajectory $u(x)$. Very importantly, the boundary values $u_{0}$ and $u_{1}$ are unknown and must be learned in-context.
The reaction-diffusion boundary-value PDE is:
$a\frac{d^{2}u(x)}{dx^{2}}+k(x)u(x)=c\,,\quad 0<x<1\,$, (3)
with $u(0)=u_{0}$ and $u(1)=u_{1}$. The forward problem is to find the solution $u(x)$ given the coefficient function $k(x)$. The inverse problem is to find the coefficient function $k(x)$ that produces the given solution $u(x)$. Here, the coefficients $a$ and $c$ and the boundary values $u_{0}$ and $u_{1}$ are all unknown and must be learned in-context.
(a)         Forward Poisson
    $\frac{d^{2}{\color[rgb]{1,0,0}G(u)(x)}}{dx^{2}}={\color[rgb]{0,0,1}u(x)}$, given $u(0),u(1)$
(b)         Inverse Poisson
    $\frac{d^{2}{\color[rgb]{0,0,1}u(x)}}{dx^{2}}={\color[rgb]{1,0,0}G(u)(x)}$
, given $u(0),u(1)$
(c)      Forward Reaction-Diffusion
   $a\frac{d^{2}}{dx^{2}}{\color[rgb]{1,0,0}u(x)}+{\color[rgb]{0,0,1}k(x)}{\color[rgb]{1,0,0}u(x)}=c$, given $a$, $c$
(d)      Inverse Reaction-Diffusion
   $a\frac{d^{2}}{dx^{2}}{\color[rgb]{0,0,1}u(x)}+{\color[rgb]{1,0,0}k(x)}{\color[rgb]{0,0,1}u(x)}=c$, given $a$, $c$
Figure 2: In-context PDE solution operator learning with the DeepOSets architecture. In each case, the top plot contains the parameter functions while the bottom plot contains the corresponding solution functions. The gray lines indicate in each case the parameter-solution function pairs given in the prompt. The blue line indicates the query parameter function, while the red line indicates the corresponding solution predicted by the trained DeepOSets model. The black line indicates the exact solution corresponding to the query, using the appropriate PDE. The DeepOsets model is trained to learn the four PDE problems simultaneously, with coefficients sampled randomly.
All PDE coefficients and boundary values are chosen randomly, so that the number of solution operators involved is in fact infinite. The physical domain in all cases is the interval $[0,1]$, which is discretized uniformly into 100 points. The prompts during both training and testing contain 4 example pairs of parameter and solution functions. The MLP $\phi$ in the DeepOSets contains 2 hidden layers of 400 neurons each, and the latent vector encoding the prompt data has size 500. The pooling operation is the sample mean. The branch and trunk networks in the DeepOSets contain hidden layers of 200, 400, and 200 neurons with intermediate layer of 150 neurons. All neural network use ReLU as the activation function.
To generate the training data, for each of the four problems (Poisson/reaction-fiffusion, forward/inverse), 100 parameter functions are sampled independently from a Gaussian process with the squared-exponential kernel (discretized on the uniform 100-point grid) with variance 2 and length scale $0.5$. For each parameter function and PDE, the boundary values $u_{0}$ and $u_{1}$ are sampled uniformly in the interval [1,1][-1,1]. In addition, in the case of the reaction-diffusion PDE, the coefficients $a$ and $c$ are sampled uniformly in the intervals $[0.025,0.075]$ and [2,2][-2,2], respectively. Then the PDE is solved for the given parameter function, boundary values, and coefficients to obtain the groundtruth solutions. The model was trained by minimizing the mean squared error loss between predicted and groundtruth solutions over a mini-batch of 100 operators sampled randomly from the training data at each iteration, using the Adam optimizer with learning rate 1e-3, for 50k iterations. Training took 14min, achieving a training loss 8.4e-3.
Sample results are shown in Figure 2. At inference time, DeepOSets can perform accurate in-context learning, without knowing which PDE, inverse or forward problem, or the value of the equation coefficients and boundary conditions are. All these unknowns are correctly disambiguated from the data in the prompt, and the solution corresponding to the query parameter is accurately predicted, without any weight updates.
6 Conclusion
We have presented DeepOSets for in-context multi-operator learning, a novel non-autoregressive, non-attention-based approach to learning solution operators of multiple parametric PDEs. By combining DeepSets for set learning with DeepONets for operator learning, DeepOSets addresses key limitations of existing approaches while maintaining strong in-context learning capabilities. The proposed architecture offers several advantages: fully parallel processing of prompts enables faster training and inference compared to sequential methods; linear complexity in prompt length provides better scalability than quadratic self-attention mechanisms; and built-in permutation-invariance provides beneficial inductive bias for training efficiency. We prove that DeepOSets are universal uniform approximators over a class of operators, which we believe is the first result of its kind to be published in the Scientific Machine Learning area. Our experiments on Poisson and reaction-diffusion problems demonstrate that DeepOSets successfully disambiguates between multiple PDEs, forward and inverse problems, and unknown equation coefficients and boundary conditions purely from prompted examples. The model achieves accurate predictions without weight updates at inference time, requiring only 14 minutes of training on all four problems simultaneously. Future work will extend DeepOSets to higher-dimensional PDEs, more diverse equation families, and longer prompts. DeepOSets establishes a computationally efficient foundation for building general-purpose PDE solvers capable of in-context adaptation, contributing toward foundation models for scientific computing.
Appendix
Let us begin by proving Lemma 4.1 showing that $(\delta,C)$-discretizations always exist.
Proof of Lemma 4.1.
Let $K\subset X$ be compact, and let $\delta>0$ be given. Choose a maximal $\delta$-separated subset of $K$, i.e., a subset $\mathcal{X}_{\delta}$ such that $d(x_{i},x_{j})\geq\delta$ for all $i\neq j$, and which is maximal with respect to this property. It is clear that the balls of radius $\delta$ around the points $x_{i}$ give a minimal covering of $K$, and so by compactness $\mathcal{X}_{\delta}$ must be a finite set. We claim that $\mathcal{X}_{\delta}$ is a $(\delta,C)$-covering for any $C>1$.
First, note that any ball of radius $\delta/2$ contains at most one point of $\mathcal{X}_{\delta}$. Indeed, fix $z\in K$ and suppose $x_{i},x_{j}\in\mathcal{X}_{\delta}$ both lie in $B(z,\delta/2)$ with $i\neq j$. Then, by the triangle inequality,
$d(x_{i},x_{j})\;\leq\;d(x_{i},z)+d(z,x_{j})\;<\;\frac{\delta}{2}+\frac{\delta}{2}\;=\;\delta,$
which contradicts the $\delta$–separation of $\mathcal{X}_{\delta}$. Hence, for every $z\in K$,
$\bigl|\{x_{i}\in\mathcal{X}_{\delta}:x_{i}\in B(z,\delta/2)\}\bigr|\leq 1.$
Next, since balls of radius $\delta$ around the points in $\mathcal{X}_{\delta}$ cover $K$, for every $y\in K$ there exists some $x_{j}\in\mathcal{X}_{\delta}$ with $d(y,x_{j})<\delta$, i.e. $x_{j}\in B(y,\delta)$. Thus, for all $y\in K$,
$\bigl|\{x_{i}\in\mathcal{X}_{\delta}:x_{i}\in B(y,\delta)\}\bigr|\geq 1.$
Combining these two observations, we obtain, for all $x,y\in K$,
$\bigl|\{x_{i}\in\mathcal{X}_{\delta}:x_{i}\in B(x,\delta/2)\}\bigr|\;<\;C\,\bigl|\{x_{i}\in\mathcal{X}_{\delta}:x_{i}\in B(y,\delta)\}\bigr|,$
for any $C>1$. Hence $\mathcal{X}_{\delta}$ is a $(\delta,C)$–discretization of $K$. ∎
Next, we state a Lemma which we will use in the proof of Theorem 4.2.
Lemma 6.1.
Let $K$ be a compact metric space and $V\subset C(K)$ be compact. Given $\varepsilon>0$, there exists $\mathbf{y}:=\{y_{1},\cdots,y_{n}\}\in K$, and $\varphi_{1},\cdots,\varphi_{n}\in C(K)$ with $\|\varphi_{i}\|\leq 1$, such that
$\|f-\sum_{i=1}^{n}f(y_{i})\varphi_{i}\|_{C(K)}<\varepsilon$
for all $f\in V$. In particular, for any $f,g\in V$ we have
$\|f-g\|_{C(K)}\leq\|f-g\|_{\ell^{\infty}(\mathbf{y})}+2\varepsilon.$ (4)
Proof.
(of Lemma 6.1) Since $V$ is compact, by the Arzela-Ascoli theorem it is also equicontinuous, and so there exists $\delta>0$ such that if $|x-y|<\delta$, then $|f(x)-f(y)|<\varepsilon$ for all $f\in V$. Since $K$ is also compact, we can pick $y_{1},\cdots,y_{n}\in K$ such that $\{B(y_{i},\delta)\}_{i=1}^{n}$ form a covering of $K$. Finally, let $\{\varphi_{i}\}_{i=1}^{n}$ be a partition of unity subordinate to the balls $\{B(y_{i},\delta)\}_{i=1}^{n}$.
Given $y\in K$, for any $f\in V$, we have that,
$|f(y)-\sum_{i=1}^{n}f(y_{i})\varphi_{i}(y)|$ $=|\sum_{i=1}^{n}(f(y)-f(y_{i}))\varphi_{i}(y)|$
$<|\sum_{i,|y-y_{i}|<\delta}(f(y)-f(y_{i}))\cdot\varphi_{i}(y)|$
$<|\sum_{i,|y-y_{i}|<\delta}\varepsilon\cdot\varphi_{i}(y))|\leq\varepsilon.$
Using this, if $f,g\in V$ we have
$\begin{split}\|f-g\|_{C(K)}&\leq\|f-\sum_{i=1}^{n}f(y_{i})\varphi_{i}\|_{C(K)}+\|f-g\|_{\ell^{\infty}(\mathbf{y})}\max_{i}\|\varphi_{i}\|_{C(K)}+\|g-\sum_{i=1}^{n}g(y_{i})\varphi_{i}\|_{C(K)}\\ &\leq\|f-g\|_{\ell^{\infty}(\mathbf{y})}+2\varepsilon.\end{split}$ (5)
We can now return to the proof of Thm 4.2.
Proof.
(of Theorem 4.2) Fix $\varepsilon>0$. Since $\mathcal{K}$ is compact and $V$ is compact, it follows that the set
$\{G(u):~G\in\mathcal{K},~u\in V\}\subset C(K_{2})$ (6)
is compact. Since $K_{2}$ is also compact, we can apply Lemma 6.1 to get that there exists a grid y of $K_{2}$, and $\{\gamma_{i}\}_{i=1}^{n}\subset C(K_{2})$, such that
$\|G(u_{q})-\sum_{i=1}^{n}G(u_{q})(y_{i})\gamma_{i}\|_{C(K_{2})}<\varepsilon/2,$
for every $G\in\mathcal{K}$ and $u_{q}\in V$. So let $\Psi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{n}$ be defined as,
$\Psi(y)=\{\gamma_{i}(y)\}_{i=1}^{n}.$
Using the notation $G(u_{q})(\textbf{y}):=(G(u_{q})(y_{1}),\cdots,G(u_{q})(y_{n}))$, we have that
$\|G(u_{q})-G(u_{q})(\textbf{y})\cdot\Psi\|_{C(K_{2})}<\varepsilon/2.$ (7)
for every $G\in\mathcal{K}$.
We now show that $G(u_{q})(\textbf{y})$ can be approximated arbitrarily well by functions of the form
$\rho\left(u_{q}(\textbf{x}),\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))\right),$ (8)
where $\textbf{x}=(x_{1},...,x_{k})$ is a grid in $K_{1}$, $\Phi$ and $\rho$ are continuous, and $u_{1},...,u_{m}$ is an arbitrary $(\delta,C)$-discretization of $V$.
Since $\mathcal{K}$ is compact, it is equi-continuous, so there exists a $\delta$ such that if $\|u-v\|_{C(K_{1})}<25\delta/6$, then $\|G(u)-G(v)\|_{C(K_{2})}<\varepsilon/(2\cdot\|\Psi\|_{C(K_{2})})$ for every $G\in\mathcal{K}$. By Lemma 6.1, there exists $\textbf{x}=\{x_{1},\cdots,x_{k}\}\in K_{1}$ such that
$\|u-v\|_{C(K_{1})}\leq\|u-v\|_{\ell^{\infty}(\mathbf{x})}+\frac{\delta}{6}.$ (9)
We let x be our grid in $K_{1}$ and $\delta$ be the value given in the Theorem.
Let $u_{1},...,u_{m}\in V$ be a $(\delta,C)$-discretization of $V$, and consider the restrictions $u_{i}(\mathbf{x})$ of $u_{i}$ to the grid $\mathbf{x}$, which lie in the compact set $V(\mathbf{x})\subset\mathbb{R}^{k}$. Let $u\in K$ and observe that by (9) we have
$\|u-u_{i}\|_{C(K_{1})}-\frac{\delta}{6}\leq\|u-u_{i}\|_{\ell^{\infty}(\mathbf{x})}\leq\|u-u_{i}\|_{C(K_{1})}$ (10)
This implies that
$\{u_{i}(\mathbf{x})\in B(u(\mathbf{x}),\delta/3)\}\subset\{u_{i}\in B(u,\delta/2)\}$ (11)
and that
$\{u_{i}\in B(u,\delta)\}\subset\{u_{i}(\mathbf{x})\in B(u(\mathbf{x}),\delta)\}.$ (12)
Hence, since $u_{1},..,u_{m}$ is a $(\delta,C)$-discretization, when restricted to the grid $\mathbf{x}$ we still have
$|\{u_{i}(\mathbf{x})\in B(u(\mathbf{x}),\delta/3)\}|\leq C|\{u_{i}(\mathbf{x})\in B(v(\mathbf{x}),\delta)\}|$ (13)
for any $u,v\in V$.
Next, we choose a maximal $2\delta$-separated subset $\widetilde{u}_{1}(\mathbf{x}),...,\widetilde{u}_{l}(\mathbf{x})$ of the set $V(\mathbf{x})\subset\mathbb{R}^{k}$, and let $\varphi_{j}:\mathbb{R}^{k}\rightarrow\mathbb{R}_{\geq 0}$ be a partition of unity subordinate to the covering $B(\widetilde{u}_{j}(\mathbf{x}),2\delta)$ of $V(\mathbf{x})$. Observe that the balls of radius $\delta$ about $\widetilde{u}_{j}(\mathbf{x})$ are disjoint, and so we have
$\varphi_{j}(v(\mathbf{x}))=1~~\text{whenever $\|v-\widetilde{u}_{j}\|_{\ell^{\infty}(\mathbf{x})}\leq\delta$.}$ (14)
In addition, by construction we have
$\varphi_{j}(v(\mathbf{x}))=0~~\text{whenever $\|v-\widetilde{u}_{j}\|_{\ell^{\infty}(\mathbf{x})}\geq 2\delta$.}$ (15)
Let $\Phi:\mathbb{R}^{k+n}\rightarrow\mathbb{R}^{l(n+1)}$ (so that $N=l(n+1)$) be defined by
$\Phi(\mathbf{u},\mathbf{v})=\{\varphi_{j}(\mathbf{u})\cdot\mathbf{v},\varphi_{j}(\mathbf{u})\}_{j=1}^{l}.$
We then have that for any $G\in\mathcal{K}$,
$\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))=\left\{\frac{1}{m}\sum_{i=1}^{m}\varphi_{j}(u_{i}(\textbf{x}))\cdot G(u_{i})(\textbf{y}),\frac{1}{m}\sum_{i=1}^{m}\varphi_{j}(u_{i}(\textbf{x}))\right\}_{j=1}^{l}$
Define $\xi:\mathbb{R}^{l(n+1)}\rightarrow\mathbb{R}^{ln}$ as,
$\xi(z_{1},\cdots,z_{l(n+1)})=\left\{\frac{(z_{j},z_{j+1},\cdots,z_{j+n-1})}{z_{j+n}}\right\}_{j\in[1,(n+1)+1,2(n+1)+1,\cdots,(l-1)(n+1)+1]}.$
Then, given the input from $\Phi$, we will have,
$\xi\left(\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))\right)=\left\{\frac{\frac{1}{m}\sum_{i=1}^{m}\varphi_{j}(u_{i}(\textbf{x}))\cdot G(u_{i})(\textbf{y})}{\frac{1}{m}\sum_{i=1}^{m}\varphi_{j}(u_{i}(\textbf{x}))}\right\}_{j=1}^{l}:=\{\overline{G(\widetilde{u}_{j})(\textbf{y})}\}_{j=1}^{l}.$
Here the $\overline{G(\widetilde{u}_{j})(\textbf{y})}$ notation is used just to emphasize that this is just the weighted average of the $G(u_{i})(\textbf{y})$ samples that are in the ball of radius $\delta$ centered at $\widetilde{u}_{j}$. We then define, $\rho:\mathbb{R}^{k+l(n+1)}\rightarrow\mathbb{R}^{n}$, as
$\rho(x,y)=\sum_{j=1}^{l}\textbf{(}\xi(y)\textbf{)}_{j}\cdot\varphi_{j}(x),$ (16)
so that we have
$\rho\left(u_{q}(\textbf{x}),\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))\right)=\sum_{j=1}^{l}\overline{G(\widetilde{u}_{j})(\textbf{y})}\cdot\varphi_{j}(u_{q}(\textbf{x})).$ (17)
Now, fix any $G\in\mathcal{K}$. We will show that the above expression approximates $G(u_{q})(\textbf{y})$. We calculate
$\left\|G(u_{q})(\textbf{y})-\sum_{j=1}^{l}\overline{G(\widetilde{u}_{j})(\textbf{y})}\cdot\varphi_{j}(u_{q}(\textbf{x}))\right\|_{\ell^{\infty}(\textbf{y})}$ $=\left\|\sum_{j=1}^{l}G(u_{q})(\textbf{y})\cdot\varphi_{j}(u_{q}(\textbf{x}))-\sum_{j=1}^{l}\overline{G(\widetilde{u}_{j})(\textbf{y})}\cdot\varphi_{j}(u_{q}(\textbf{x}))\right\|_{\ell^{\infty}(\textbf{y})}$
$=\left\|\sum_{j,\|u_{q}-\widetilde{u_{j}}\|_{\ell^{\infty}(\textbf{x})}<2\delta}(G(u_{q})(\textbf{y})-\overline{G(\widetilde{u}_{j})(\textbf{y}))}\cdot\varphi_{j}(u_{q}(\textbf{x}))\right\|_{\ell^{\infty}(\textbf{y})}$
Then, note
$\left\|G(u_{q})(\textbf{y})-\overline{G(\widetilde{u}_{j})(\textbf{y})}\right\|_{\ell^{\infty}(\textbf{y})}$ $=\left\|G(u_{q})(\textbf{y})-\frac{\sum_{i=1}^{m}\varphi_{j}(u_{i}(\textbf{x}))\cdot G(u_{i})(\textbf{y})}{\sum_{i=1}^{m}\varphi_{j}(u_{i}(\textbf{x}))}\right\|_{\ell^{\infty}(\textbf{y})}$
$=\left\|G(u_{q})(\textbf{y})-\sum_{i=1}^{m}\frac{\varphi_{j}(u_{i}(\textbf{x}))}{\sum_{i=1}^{m}\varphi_{j}(u_{i}(\textbf{x}))}\cdot G(u_{i})(\textbf{y})\right\|_{\ell^{\infty}(\textbf{y})}$
$=\left\|\sum_{i=1}^{m}\frac{\varphi_{j}(u_{i}(\textbf{x}))}{\sum_{i=1}^{m}\varphi_{j}(u_{i}(\textbf{x}))}\cdot\left(G(u_{q})(\textbf{y})-G(u_{i})(\textbf{y})\right)\right\|_{\ell^{\infty}(\textbf{y})}$
$\leq\varepsilon/(2\cdot\|\Psi\|_{C(K_{2})})\left|\sum_{i=1}^{m}\frac{\varphi_{j}(u_{i}(\textbf{x}))}{\sum_{i=1}^{m}\varphi_{j}(u_{i}(\textbf{x}))}\right|=\varepsilon/(2\cdot\|\Psi\|_{C(K_{2})})$
Where we use that $\|u_{q}-\widetilde{u}_{j}\|_{\ell^{\infty}(\textbf{x})}<2\delta$, $\|\widetilde{u}_{j}-u_{i}\|_{\ell^{\infty}(\textbf{x})}<2\delta$, and (9) imply that
$\|u_{q}-u_{i}\|_{C(K_{1})}\leq\|u_{q}-\widetilde{u}_{j}\|_{\ell^{\infty}(\textbf{x})}+\|\widetilde{u}_{j}-u_{i}\|_{\ell^{\infty}(\textbf{x})}+\delta/6<25\delta/6,$ (18)
so that by the choice of $\delta$, $\|G(u_{q})-G(u_{i})\|_{C(K_{2})}<\varepsilon/(2\cdot\|\Psi\|_{C(K_{2})})$.
So we have
$\left\|G(u_{q})(\textbf{y})-\sum_{j=1}^{l}\overline{G(\widetilde{u}_{j})(\textbf{y})}\cdot\varphi_{j}(u_{q}(\textbf{x}))\right\|_{\ell^{\infty}(\textbf{y})}\leq\varepsilon/(2\cdot\|\Psi\|_{C(K_{2})}),$
which combining with (7) and using (17) gives us,
$\|G(u_{q})-\rho\left(u_{q}(\textbf{x}),\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))\right)\cdot\Psi\|_{C(K)}$
$\leq\|G(u_{q})-G(u_{q})(\textbf{y})\cdot\Psi\|_{C(K)}+\|G(u_{q})(\textbf{y})\cdot\Psi-\rho\left(u_{q}(\textbf{x}),\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))\right)\cdot\Psi\|_{C(K)}$
$<\varepsilon/2+\varepsilon/2=\varepsilon.$
Finally, we need to show that $\rho$ and $\Phi$ can be chosen continuous. It is clear from the definition that $\Phi$ is continuous. To show that $\rho$ is also continuous, we need to show that the denominator
$\frac{1}{m}\sum_{i=1}^{m}\varphi_{j}(u_{i}(\textbf{x}))$ (19)
in the definition of $\xi$ is bounded away from zero uniformly for any $(\delta,C)$-discretization $u_{1},...,u_{m}$ of $V$.
To prove this, we use (13) and (14). Let $B(v_{1}(\textbf{x}),\delta/3),...,B(v_{t}(\textbf{x}),\delta/3)$ be a covering of $V(\textbf{x})$. It follows that one of these balls must contain at least $m/t$ elements $u_{i}(\textbf{x})$ (recall that $m$ is the size of the $(\delta,C)$-discretization, which is not controlled). Using (13) this means each ball $B(u_{i}(\textbf{x}),\delta)$ must contain at least $m/(Ct)$ elements of the $(\delta,C)$-discretization. Finally, using (14) this gives
$\frac{1}{m}\sum_{i=1}^{m}\varphi_{j}(u_{i}(\textbf{x}))\geq\frac{1}{m}\frac{m}{Ct}=\frac{1}{Ct}>0.$ (20)
This latter quantity is bounded independently of the $(\delta,C)$-discretization. This completes the proof. ∎